# Implementation Study of Cost-Effective Verification for Pietrzak’s VDF in Ethereum Smart Contracts

### v4

@@ 4.4 Halving Cost
The results of our experiments demonstrate the anticipated correlation. Figure 2(a) and Figure 2(b) display the gas cost of the repeated halving operations with $\lambda=2048$ and $\lambda=3072$ correspondingly. The data exhibits a clear linear relationship, allowing us to perform a regression analysis and model the results using a linear function. By employing the method of least squares approximation, we obtain the subsequent functions as Equation 5 and 6:
@@ 5.1 Theoretical Analysis on Proof Generation for Gas Usage Optimization
The optimized proof generation in this research means minimizing the verification gas cost in EVM. We can define the total gas cost function $g_{total}$ using the previous regression results.
$\mathcal{G}_{total}(\tau,\delta)=\mathcal{G}_{calldata}+\mathcal{G}_{halving}+\mathcal{G}_{exp}+C$ (7)
$=\big{(}\alpha\cdot(\tau-\delta)+c_{1}\big{)}+\big{(}\beta\cdot(\tau-\delta)+c_{2}\big{)}+\big{(}\gamma\cdot 2^{\delta}+c_{3}\big{)}+C$ (8)
@@
From Theorem 1, we can get the unique $\delta_{m}$ optimizing the gas cost. Applying all the regression results in Equation 1, 2, 3, 4, 5, and 6, we get the optimizing $\delta_{m}\approx 9.43$ for $\lambda=2048$ and $\delta_{m}\approx 9.15$ for $\lambda=3072$ respectively. In practice, $\delta$ for implementation must be an integer. Therefore, we can get the practical integer $\delta$ with the following corollary.
@@ 5.2 Optimized Verification Gas Cost
Figures 4(a) and 4(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively. Consistent with the theoretical predictions in Theorem 1, the gas usage graph shows a decline to a minimal point before increasing sharply. Optimal gas costs were observed at $\delta=9$ for both configurations, aligning perfectly with the findings of Corollary 2. This alignment between theoretical expectations and empirical results confirms the accuracy of our model and highlights its potential for optimizing Pietrzak VDF verification in blockchain applications.
@@
6 Conlusion and Future Works
Our study demonstrates that Pietrzak’s VDF can be effectively optimized for the EVM, reducing gas costs from 4M to 2M and minimizing proof size to under 8 KB with a 2048-bit RSA key. These optimizations validate the theoretical models and show practical viability. While our paper does not address the Wesolowski VDF implementation, which poses challenges due to its hash-to-prime function, exploring indirect or approximate implementation methods could make it feasible for the EVM. Future work should focus on utilizing Wesolowski VDF verification on EVM as it has a big benefit on the constant and short proof length and verification cost. We plan to implement on-chain Wesolowski VDF verification both directly and indirectly. For indirect implementation, a hash-to-prime function could be executed in a Truebit-style proof-of-fraud system. For direct implementation, refining the hash-to-prime function to be viable and secure on the EVM is essential.
@@ Appendix A Code availability
Appendix B Proof of Theorem 1
@@
Appendix C Proof of Corollary 2

### v5

@@ 4.4 Halving Cost
The results of our experiments demonstrate the anticipated correlation. Figure 2(c) and Figure 2(d) display the gas cost of the repeated halving operations with $\lambda=2048$ and $\lambda=3072$ correspondingly. The data exhibits a clear linear relationship, allowing us to perform a regression analysis and model the results using a linear function. By employing the method of least squares approximation, we obtain the subsequent functions as Equation 6 and 7:
@@ 5.1 Theoretical Analysis on Proof Generation for Gas Usage Optimization
The optimized proof generation in this research means minimizing the verification gas cost in EVM. We can define the total gas cost function $g_{total}$ using the regression results (Eq. 27) in the previous section.
$\mathcal{G}_{\text{total}}(\tau,\delta)=\mathcal{G}_{\text{calldata}}+\mathcal{G}_{\text{halving}}+\mathcal{G}_{\text{exp}}+C$ (8)
$=\big{(}\underbrace{\alpha\cdot(\tau-\delta)+c_{1}}_{\mathcal{G}_{\text{calldata}}}\big{)}+\big{(}\underbrace{\beta\cdot(\tau-\delta)+c_{2}}_{\mathcal{G}_{\text{halving}}}\big{)}+\big{(}\underbrace{\gamma\cdot 2^{\delta}+c_{3}}_{\mathcal{G}_{\text{exp}}}\big{)}+C$ (9)
@@
From Theorem 1, we can get the unique $\delta_{m}$ optimizing the gas cost. Applying all the regression results in Equation 2, 3, 4, 5, 6, and 7, we get the optimizing $\delta_{m}\approx 9.43$ for $\lambda=2048$ and $\delta_{m}\approx 9.15$ for $\lambda=3072$ respectively. In practice, $\delta$ for implementation must be an integer. Therefore, we can get the practical integer $\delta$ with the following corollary.
@@ 5.2 Optimized Verification Gas Cost
Figures 3(a) and 3(b) depict the total gas cost for Pietrzak VDF verification at $\lambda=2048$ and $\lambda=3072$ respectively. Consistent with the theoretical predictions in Theorem 1, the gas usage graph shows a decline to a minimal point before increasing sharply. Optimal gas costs were observed at $\delta=9$ for both configurations, aligning perfectly with the findings of Corollary 2. This alignment between theoretical expectations and empirical results confirms the accuracy of our model and highlights its potential for optimizing Pietrzak VDF verification in blockchain applications.
@@
6 Discussion
This section examines the broader implications and feasibility of implementing Pietrzak VDF verification on Ethereum. First, we consider the feasibility of the observed gas costs and discuss how frequent verification might remain economically viable on both Ethereum’s Layer 1 and Layer 2. Next, we address the lack of an established economic model for VDFs and the challenges it poses for incentivizing participants in decentralized environments. Lastly, we look beyond the EVM and outline how the findings could extend to more general computing contexts, suggesting avenues for future research and optimization.
6.1 Practical Feasibility of Our Observed Costs.
From Section 5, we note that even with an optimal parameter of $\delta=9$, the gas cost for verifying a Pietrzak VDF can reach a few million gas units. On Ethereum Layer 1, this is still somewhat expensive but may be acceptable for security-critical operations that are not executed frequently (e.g., infrequent random-beacon generation or governance steps). However, in Layer 2 environments with significantly reduced gas prices, these same verification procedures can become much more practical and cost-effective, especially for repeated or more frequent VDF verifications. By striking this balance—an optimal $\delta$ and a cheaper execution environment—our approach could feasibly support advanced use cases on EVM-compatible rollups, offering robust cryptographic guarantees at a fraction of Layer 1 costs.
6.2 Lack of Economic VDF Protocols in Ethereum
Integrating VDFs into smart contracts offers significant potential to enhance blockchain security and reliability, yet faces major challenges due to the lack of a robust economic model tailored to their unique operational needs, such as incentivizing both provers and verifiers and managing commit-reveal-recover protocols. This absence hampers practical deployment, as it fails to compensate those who contribute significant computational resources, deterring potential contributors and affecting scalability and effectiveness in real-world applications. Additionally, the decentralized nature of blockchain necessitates economic strategies that support technical operations and ensure sustainability without central governance, promoting honest participation and deterring dishonesty to maintain integrity and fairness within the blockchain ecosystem. Thus, innovative economic models specifically designed for decentralized environments are crucial to make VDFs both technically effective and economically viable.
6.3 Future Works on a General Environment
Through this research, we investigated the efficient implementation of Pietrzak VDF verification within the Ethereum Virtual Machine (EVM) environment and observed results that align both theoretically and experimentally. During our study, we identified significant effects of generalized proof generation and verification, which were also included within the predictable realms of theoretical analysis. We anticipate that such an approach could be applicable not only within the EVM but also in more general computing environments. Although resource pricing in general systems would differ from the gas costs in the EVM, with elements like the repetition of halving protocols and resource consumption proportional to the proof length being predictable, the costs of exponentiation in VDF ($2^{2^{T}}$) are expected to be exponential. Therefore, based on the outcomes of this study, further research on optimizing Pietrzak VDFs in general computing environments appears reasonable.
7 Conclusion and Future Works
Our study demonstrates that Pietrzak’s VDF can be effectively optimized for the EVM, reducing gas costs from 4M to 2M and minimizing proof size to under 8 KB with a 2048-bit RSA key. These optimizations validate the theoretical models and show practical viability. While our paper does not address the Wesolowski VDF implementation, which poses challenges due to its hash-to-prime function, exploring indirect or approximate implementation methods could make it feasible for the EVM. Future work should focus on utilizing Wesolowski VDF verification on EVM as it has a big benefit on the constant and short proof length and verification cost. We plan to implement on-chain Wesolowski VDF verification both directly and indirectly. For indirect implementation, we suggest executing the hash-to-prime function in an off-chain, challenge-based verification system (inspired by Truebit [^4] Truebit [30] is a protocol that outsources computation off-chain while allowing anyone to challenge an allegedly correct result within a specific dispute window. If the challenge proves the result incorrect, a fraud proof is submitted on-chain, ensuring only correct computations are ultimately accepted.). This approach enables a “proof-of-fraud” mechanism, where a claimed correct computation can be disputed during a challenge period, thereby minimizing on-chain costs without compromising correctness.
@@ Appendix A Code availability
Appendix B Expected Gas Cost for Hash-to-Prime in Smart Contracts
In Wesolowski’s VDF, implementing the hash-to-prime verification on-chain requires an algorithm capable of determining whether numerous candidate values are prime. As efficient probabilistic methods for prime checking, we considered Miller–Rabin primality test with 11 iterations, and the Baillie–PSW primality test which uses a single iteration of Miller–Rabin to base 2 and a Lucas primality test [31]. For evaluation, we tested 6,100 randomly selected primes over $2^{255}$. On average, verifying a single prime consumed 21,446 gas for Miller–Rabin test and 45,912 gas for Baillie–PSW test. All prime number test cases, test scripts, and results are available in the repository referenced in A.
In the worst case, the number of prime candidates can approach $\bigl{(}\log x\bigr{)}^{2}$ for large $x$, following Shanks’s upper bound on gaps between consecutive primes [32]. Since we are dealing with the maximum number $12^{256}-$ 1, we set $x=2^{256}$, giving $\bigl{(}\log(2^{256})\bigr{)}^{2}\approx 5939$. Verifying 5939 candidate primes, each requiring about 21,446 gas for Miller–Rabin test or 45,912 gas for Baillie–PSW test, would quickly exceed the gas limit and render a purely on-chain implementation impractical. Moreover, even a single Miller–Rabin check can cost about \$0.28 (at 5 gwei) to \$1.12 (at 20 gwei) with ETH at \$2600, so multiplying by 5939 quickly runs into thousands of dollars, further underscoring the infeasibility of on-chain hash-to-prime.
Appendix C Proof of Theorem 1
@@
Appendix D Proof of Corollary 2